# Flow Graph-Based Classification of Defects4J Faults

### v1

@@ Abstract
Software fault datasets such as Defects4J provide for each individual fault its location and repair, but do not characterize the faults. Current classifications use the repairs as proxies, which does not capture the intrinsic nature of the fault. In this paper, we propose a new, direct fault classification scheme based on the control- and data-flow graph representations of the program. Our scheme comprises six control-flow and two data-flow fault classes. We apply this to 488 faults from seven projects in the Defects4J dataset. We find that one of the data-flow fault classes (definition fault) is the most common individual class but that the majority of faults are classified with at least one control-flow fault class. The majority of the faults are assigned between one and three classes. Our proposed classification can be applied to other fault datasets and can be used to improve fault localization and automated program repair techniques for specific fault classes.
I Introduction
@@
Software debugging is an expensive and time-consuming process [1, 2], requiring faults to be found (fault localization) and fixed (fault repair). Its effectiveness depends on several factors, but one factor that is often overlooked is the nature of the fault itself [3]. Typically, this fault nature is unknown, but an analysis of known real-world faults allows us to identify common fault types, which can in turn be used to improve fault localization and repair techniques.
@@
However, investigating the intrinsic nature of faults can be labor-intensive, and classifications typically use the applied repairs as proxies [4, 3, 5, 2]. This allows automating the classification process but it does not necessarily capture the intrinsic nature of the faults themselves, since faults can typically be repaired in different ways.
@@
In this paper, we describe a new fault classification scheme based on control- and data-flow graphs. The main novelty of our scheme is its grounding in the (abstract) structural properties of the faults themselves, rather than the applied repairs. This not only eliminates any potential bias by repair-based classifications, it also emphasizes the commonalities of different faults under syntactic variations. Consider for example the faults (and repairs) shown in LABEL:listing:lang62 and LABEL:listing:chart18. In both cases, the underlying problem is that the control flow from the fault location to the next location is wrong (i.e., whenever the program executes the faulty statement, it executes the wrong statement next), despite the differences in the faulty statements (i.e., return vs. missing statement) and the repair (i.e., substitution of return by throw vs. insertion of break). We thus classify both faults as (unconditional) control-flow target faults.
@@
Our scheme comprises six control-flow and two data-flow fault classes. We use it to classify 488 faults from seven projects in the Defects4J dataset; the classification results are available at https://github.com/alexvdspuy/fcd4j. The most common individual class is one of the data-flow fault classes (definition fault); however, the control-flow classes are overall more common. Similar to repair classes [2], our fault classes are not disjoint since a single fault can have characteristics of multiple classes. The majority of the faults are thus assigned one to three of the eight classes.
II Flow Graph-Based Classification
II-A Flow graphs
@@
We base our classification on flow graph representations of programs. Specifically, we consider the control-flow graph (CFG) and data-flow graph (DFG) as a single structure $G=(V,E_{C},E_{D})$, where the vertices $V$ correspond to the program statements, the CFG edges $E_{C}$ correspond to changes in the program counter, and the DFG edges $E_{D}$ represent associations between variable writes or defs and their references or uses. The flow graph construction, with explicit def-use association edges, is fairly standard [6, 7]. For the classification we assume that ($i$) jumps (i.e., break-, continue-, return-, and throw-statements) are represented by CFG edges only, and not by vertices, and ($ii$) method calls are represented by a vertex modeling the parameter passing and a CFG edge modeling the jump. We list our fault classes in §II-B to II-C.
II-B Control-flow faults
II-B1 Unconditional control-flow target faults
@@
Such faults occur when at least one CFG edge has the wrong target and the source of the faulty edge has no other outgoing edge, i.e., is not a predicate node. Any run that reaches the source will thus traverse the faulty edge and inevitably enter an error state, but not necessarily fail. Faulty edges can correspond to wrong statement orders, or to missing, extraneous, or wrong jumps and method calls. Furthermore, the flow graph must already have the correct set of vertices, unless the fault occurs in combination with other faults.
@@
Such faults occur when a sequence of consecutive edges connect multiple vertices in the wrong order, but the set of source and target vertices of these edges is correct; a repair rearranges the edges’ sources and targets, which corresponds to reordering the statements. Closure-$13$ shows an example:
@@
Listing 1: Statement order fault (Closure-$13$)
@@
Listing 2: Jump fault—wrong jump target (Chart-$18$)
In contrast, Lang-$62$ contains a missing jump fault where the control flow incorrectly connects the assignment to the next case-label, and the insertion of the break-statement repairs this. Similarly, removing a faulty extraneous jump restores the normal control flow.
@@
Listing 3: Jump fault—missing jump (Lang-$62$)
@@
The similarity between jumps and method calls also extends to faults. The only small difference is that the repair of a missing method call with arguments also requires inserting the vertex that models the parameter passing; this implies that such faults are also definition faults (§II-C1). Listings LABEL:listing:time22 and LABEL:listing:math64 show a wrong resp. missing method call (cf. Listings LABEL:listing:chart18 and LABEL:listing:lang62 for the corresponding jump faults).
@@
Listing 4: Method call fault—wrong method (Time-$22$)
@@
Listing 5: Method call fault—missing call (Math-$64$)
II-B2 Control-flow predicate faults
Control-flow predicates (i.e., the conditions of conditionals and loops) determine which CFG paths a program follows. A control-flow predicate fault occurs if at least one of these paths is incorrect. This can be caused by a fault in the condition itself, missing/extraneous predicate vertices guarding otherwise correct existing vertices, or entire predicate blocks that are missing.
@@ II-B2 Control-flow predicate faults
Predicate vertex faults
This class captures faults where the CFG structure is correct, but the label of a predicate vertex (i.e., the condition) is wrong. For example, Chart-$1$ contains a predicate vertex fault, but if the faulty vertex is also the target of a faulty DFG edge (i.e., references a wrong variable), it is also classified as a data-flow fault.
@@
Listing 6: Predicate vertex fault (Chart-$1$)
@@
This class captures faults where the CFG has an extraneous (resp. missing) predicate node guarding existing vertices, leading to paths that execute the guarded vertices in too few (resp. many) paths. For example, Lang-$55$ contains a missing if-guard fault:
@@
Listing 7: Predicate existence fault—missing if (Lang-$55$)
@@
In contrast to predicate existence faults, predicate block faults capture faults where the CFG is missing not only the predicate vertex but also the entire guarded block. Closure-$85$ contains such a fault:
@@
Listing 8: Predicate block fault—missing while (Closure-$85$)
II-C Data-flow faults
Data-flow faults occur when the value written into a memory location does not match the requirements of a subsequent read from that location. This can be caused by faulty or missing definitions, or by faulty uses.
II-C1 Definition faults
Such faults occur at the source of a DFG edge. We distinguish between a wrong write value (i.e., the wrong value is written to the correct variable, see LABEL:listing:closure97), wrong write location (i.e., the correct value is written to the wrong variable), or a missing definition.
@@ II-C1 Definition faults
Listing 9: Definition fault—wrong write value (Closure-$97$)
II-C2 Use faults
Such faults occur at the target of a DFG edge, where the variable is read from the wrong location. Jsoup-57 contains a faulty use of attributes instead of it:
@@ II-C2 Use faults
Listing 10: Use fault (Jsoup-$57$)
III Dataset Construction Methodology
We use Defects4J [8], a repository of real-world Java projects, as our data source. We classify the 488 faults from seven Defects4J projects (see Section III) with manual thematic analysis, similar to the process described by Sobreira et al. [2]. Each fault is inspected in terms of the source and repaired program, as well as failing test cases. The fault is then assigned the appropriate fault classes from our classification scheme. The first author of this paper performed manual classification, which was then reviewed by the second author, and another round of classification was done.
TABLE I: Project characteristics [7, 8, 9] and fault class frequency (maximum per project bolded)
{tblr}
@@ III Dataset Construction Methodology
Project KLOC order jump call pred guard block def use Avg per fault CF DF CF/DF All
[…]
Closure $90$ $6$ $9$ $37$ 4242{42}42 $53$ $30$ $\mathbf{81}$ $6$ $1.98\pm 1.03$ $52$ $14$ $67$ $133$
Lang $22$ $2$ $18$ $11$ $11$ $31$ $11$ $\mathbf{43}$ $8$ $2.08\pm 0.86$ $20$ $12$ $33$ $65$
Math $85$ $1$ $17$ $29$ $24$ $36$ $18$ $\mathbf{83}$ $20$ $2.15\pm 1.04$ $22$ $27$ $57$ $106$
Mockito $11$ 00 $3$ $15$ $2$ $16$ $6$ $\mathbf{28}$ $4$ $1.95\pm 0.86$ $10$ $4$ $24$ $38$
Time $28$ $3$ $9$ $6$ 77{7}7 $17$ $3$ $\mathbf{21}$ $5$ $2.63\pm 1.61$ $6$ $4$ $17$ $27$
[dotted] Jsoup $10$ $3$ $4$ $36$ $14$ $30$ $23$ $\mathbf{59}$ $6$ $1.88\pm 0.85$ $33$ $14$ $46$ $93$
[…]
@@
We consider the difference in the flow graphs of the faulty and fixed code to manually assign fault classes according to the intrinsic fault nature. We infer the nature of the fault from the program context, taking into account the repair semantics. However, we investigate alternative repairs and compare all repair flow graphs to the faulty flow graph—the applied repair does not drive our classification.
We note that Sobreira et al. [2] were unable to classify 22 of the 395 faults in terms of their repair patterns; we are able to classify these faults in terms of our classification scheme. We also consider the Jsoup project, which has not been classified before and as such is unbiased by other classifications.
III-A Limitations
@@
We consider, similarly to Sobreira et al. [2], each Defects4J fault entry as a single fault that may display characteristics of multiple classes. However, certain entries do contain multiple independent faults (e.g., Chart-$18$). For comparability and compatibility with the repair-based classification [2], we do not distinguish between multi-faults within an entry.
@@
Several of the Defects4J patches combine the repair with refactoring (see LABEL:listing:math46). This biases a repair-based classification and may inhibit our fault classification.
IV Dataset Statistics

### v2

# An Anatomy of 488 Faults from Defects4J Based on the Control- and Data-Flow Graph Representations of Programs

@@ Abstract.
Software fault datasets such as Defects4J provide for each individual fault its location and repair, but do not characterize the faults. Current classifications use the repairs as proxies, but these do not capture the intrinsic nature of the fault. In this paper, we propose a new, direct fault classification scheme based on the control- and data-flow graph representations of programs. Our scheme comprises six control-flow and two data-flow fault classes. We manually apply this scheme to 488 faults from seven projects in the Defects4J dataset. We find that the majority of the faults are assigned between one and three classes. We also find that one of the data-flow fault classes (definition fault) is the most common individual class but that the majority of faults are classified with at least one control-flow fault class. Our proposed classification can be applied to other fault datasets and can be used to improve fault localization and automated program repair techniques for specific fault classes.
software faults, fault classification, taxonomy
conference: The 29th International Conference on Evaluation and Assessment in Software Engineering; 17–20 June, 2025; Istanbul, Turkeyccs: Software and its engineering Software defect analysiscopyright: none
1. Introduction
@@
Software debugging is an expensive and time-consuming process (DiGiuseppe and Jones, 2015; Sobreira et al., 2018), requiring faults to be found (fault localization) and fixed (fault repair). Its effectiveness depends on several factors, but one factor that is often overlooked is the nature of the fault itself (Nath et al., 2012). Typically, this fault nature is unknown, but an analysis of known real-world faults allows us to identify common fault types, which can in turn be used to improve fault localization and repair techniques.
@@
However, investigating the intrinsic nature of faults can be labor-intensive, and classifications typically use the applied repairs as proxies (Pan et al., 2009; Nath et al., 2012; Lucia et al., 2014; Sobreira et al., 2018). This allows automating the classification process but it does not necessarily capture the intrinsic nature of the faults themselves, since faults can typically be repaired in different ways. Consider for example LABEL:listing:lang7 and LABEL:listing:math46:
- if (str.startsWith("--")) {
- return null;
Listing 1: Conditional block repair—removal (Lang-7) (Sobreira et al., 2018)
In LABEL:listing:lang7, the underlying fault is not the conditional, but the return-statement; removing only the return would also be a valid fix.
- return isZero ? NaN : INF;
+ return NaN;
Listing 2: Unwraps-from repair—if-else (Math-46) (Sobreira et al., 2018)
For LABEL:listing:math46, the repair is classified as unwraps-from because the conditional is removed, but the true fault is in referencing and returning INF instead of NaN; another possible repair is to keep the conditional and simply replace INF with NaN.
@@
In this paper, we describe a new fault classification scheme based on the control- and data-flow graph representations of programs, and a corresponding dataset. The main novelty of our scheme is its grounding in the (abstract) structural properties of the faults themselves, rather than the applied repairs. This emphasizes the commonalities of different faults under syntactic variations. Consider for example the faults (and repairs) shown in LABEL:listing:lang62 and LABEL:listing:chart18. In both cases, the underlying problem is that the control flow from the fault location to the next location is wrong (i.e., whenever the program executes the faulty statement, it executes the wrong statement next), despite the differences in the faulty statements (i.e., return vs. missing statement) and the repair (i.e., substitution of return by throw vs. insertion of break). We thus classify both faults as unconditional control-flow target faults.
@@
Our scheme comprises six control-flow and two data-flow fault classes. We manually classify 488 faults from seven projects in the Defects4J dataset accordingly. The most common individual class is one of the data-flow fault classes (definition faults); however, the control-flow classes are overall more common. Similar to repair classes (Sobreira et al., 2018), our fault classes are not disjoint since a single fault can have characteristics of multiple classes. The majority of the faults are thus assigned one to three of the eight classes.
Our classification is useful for investigating intrinsic characteristics of faults, rather than their repairs. We currently use it to evaluate the effectiveness of different spectrum-based fault localization (SBFL) techniques over the different fault classes, and our preliminary results show that SBFL is more effective for pure data-flow than pure control-flow faults. We also see possible applications in automated program repair. For example, we can use the fault class frequency information to guide automated program repair to attempt fixes in order of fault class frequency; if we can classify individual faults on-the-fly, we could even tailor individual repair operations to the specific fault classes.
In this paper, we make two main contributions. First, we introduce a novel flow graph-based fault classification scheme, which we present with examples in Section 2. Second, we construct a dataset of faults classified according to this scheme. We explain our construction methodology in Section 3 and analyze the dataset statistics in Section 4; dataset and replication package are publicly available (see Section 5). We then consider several ways to use the dataset to answer research questions in Section 6. We summarize related work in Section 7 and our conclusions in Section 8.
2. Flow Graph-Based Classification
2.1. Flow graphs
@@
We base our classification on flow graph representations of programs. Specifically, we consider the control-flow graph (CFG) and data-flow graph (DFG) as a single structure $G=(V,E_{C},E_{D})$, where the nodes $V$ correspond to the program statements, the CFG edges $E_{C}$ correspond to changes in the program counter, and the DFG edges $E_{D}$ represent associations between variable writes or defs and their references or uses. The flow graph construction, with explicit def-use association edges, is fairly standard (Allen, 1970; Ribeiro et al., 2019). For the classification we assume that ($i$) jumps (i.e., break-, continue-, return-, and throw-statements) are represented by CFG edges only, and not by nodes, and ($ii$) method calls are represented by a node modeling the parameter passing and a CFG edge modeling the jump.
2.2. Control-flow faults
2.2.1. Unconditional control-flow target faults
@@
Such faults occur when an CFG edge has the wrong target and the source of this faulty edge has no other outgoing edge, i.e., does not leave from a predicate node. Any run that reaches the source node will thus traverse the faulty edge and inevitably enter an error state, but not necessarily fail. Faulty edges can correspond to wrong statement orders, or to missing, extraneous, or wrong jumps and method calls; we thus distinguish three different classes of unconditional control-flow faults, as discussed below. Furthermore, for classification as an unconditional control-flow fault we require that the flow graph already contains the correct set of nodes, unless the fault occurs in combination with other faults.
@@
Such faults occur when a sequence of consecutive edges connect multiple nodes in the wrong order, but the set of source and target nodes of these edges is correct; a repair rearranges the edges’ sources and targets, which corresponds at the source level to a reordering of the statements. Closure-13 shows an example:
@@
Listing 3: Statement order fault (Closure-13)
@@
Listing 4: Jump fault—wrong jump target (Chart-18)
In contrast, Lang-62 contains a missing jump fault where the control flow incorrectly connects the assignment to the next case-label, and the insertion of the break-statement repairs this. Conversely, removing a faulty extraneous jump restores the normal control flow.
@@
Listing 5: Jump fault—missing jump (Lang-62)
@@
The similarity between jumps and method calls also extends from the CFG structure to faults. The only small difference is that the repair of a missing method call with arguments also requires inserting the node that models the parameter passing; this implies that such faults are also definition faults (§2.3.1). Listings LABEL:listing:time22 and LABEL:listing:math64 show a wrong resp. missing method call (cf. Listings LABEL:listing:chart18 and LABEL:listing:lang62 for the corresponding jump faults).
@@
Listing 6: Method call fault—wrong method (Time-22)
@@
Listing 7: Method call fault—missing call (Math-64)
2.2.2. Control-flow predicate faults
Control-flow predicates (i.e., the conditions of conditionals and loops) determine which CFG paths a program follows. A control-flow predicate fault occurs if at least one of these paths is incorrect. This can be caused by a fault in the condition itself, missing/extraneous predicate nodes guarding otherwise correct existing nodes, or entire predicate blocks that are missing or superfluous. We therefore distinguish three different classes of control-flow predicate faults, as discussed below.
@@ 2.2.2. Control-flow predicate faults
Predicate node faults
This class captures faults where the CFG structure is correct, but the label of a predicate node (i.e., the condition) is wrong. For example, Chart-1 contains a predicate node fault, but if the faulty node is also the target of a faulty DFG edge (i.e., references a wrong variable), it is additionally classified as a data-flow fault.
@@
Listing 8: Predicate node fault (Chart-1)
@@
This class captures faults where the CFG has an extraneous (resp. missing) predicate node guarding existing nodes, leading to paths that execute the guarded nodes in too few (resp. many) paths. For example, Lang-55 contains a missing if-guard fault, which causes the existing assignment to stopTime to be executed too often:
@@
Listing 9: Predicate existence fault—missing if (Lang-55)
@@
In contrast to predicate existence faults, predicate block faults capture situations where the CFG is missing not only the predicate node but also the entire guarded block. Closure-85 exhibits such a fault:
@@
Listing 10: Predicate block fault—missing while (Closure-85)
2.3. Data-flow faults
Data-flow faults occur when the value transported along a DFG edge does not match the requirements of the read at the target of this edge. The edge itself can be correct but the corresponding value is incorrect, or the source of the edge can be wrong, either caused by a faulty or a missing definition, or by a faulty use.
2.3.1. Definition faults
Such faults occur at the source of a DFG edge. In the simplest case, the source location is correct, but the value written to the memory location (i.e., variable) used at the target location is wrong. Closure-97 exhibits such a fault; note that this is not a use fault, despite the variable change from lvalInt to lvalLong, because lvalLong does not exist in the faulty program.
@@ 2.3.1. Definition faults
Listing 11: Definition fault—wrong write value (Closure-97)
Alternatively, the source location of the edge may be wrong, either because the correct value is written to the wrong location (see LABEL:listing:math21), or because a required write is missing (see LABEL:listing:lang17).
// find maximal diagonal element
- swap[r] = r;
+ int swapR = r;
for (int i = r + 1; i < order; ++i) {
Listing 12: Definition fault—wrong location (Math-21)
out.write(c);
+ pos+= c.length;
Listing 13: Definition fault—missing definition (Lang-17)
2.3.2. Use faults
Such faults occur at the target of a DFG edge, where the value is read from the wrong location. Jsoup-57 contains a faulty use of attributes instead of it:
@@ 2.3.2. Use faults
Listing 14: Use fault (Jsoup-57)
Note that both wrong location definition faults and use faults involve a wrong DFG edge source location; the difference is determined by the corresponding variable: if the variable read is correct, the left-hand side of an assignment needs to be changed to “rewire” the edge, otherwise the used variable needs to be changed. Note also that use faults often occur in conjunction with other faults, e.g., a definition or method call fault. The latter occurs in LABEL:listing:jsoup57 because remove(attrKey) is called instead of remove().
3. Dataset Construction Methodology
We use Defects4J (Just et al., 2014), a repository of real-world Java projects, as our data source. Defects4J is often employed to test the effectiveness of fault localization and repair techniques (Marcilio, 2023; Pearson et al., 2017; Sobreira et al., 2018), making it suited for our aim of creating a classification dataset for testing purposes. We classify all 488 faults from seven Defects4J projects (see Table 1) with manual thematic analysis (Sobreira et al., 2018). Our classification can be compared to Sobreira et al. (Sobreira et al., 2018)’s repair classification of 395 of these faults, thus we use Defects4J v1.2.0 (Def, 2018) for the relevant projects. However, we consider another 93 faults of Jsoup, heretofore unclassified, in Defects4J v3.0.1(Def, 2024).
Each fault is inspected in terms of the source and repaired program, as well as failing test cases. The fault is then assigned the appropriate fault classes from our classification scheme. The first author of this paper performed manual classification, which was validated by the second author’s review, and the classification was adjusted accordingly. Disagreements regarding the classification of individual faults were resolved via discussions between both authors. Manual classification was used to gain an understanding of the intrinsic nature of each fault, allowing us to take into account subtleties that an algorithm may overlook.
Table 1. Project characteristics (Ribeiro et al., 2019; Just et al., 2014; Marcilio, 2023), frequency of fault class assignments, and frequency of distinct fault types
Fault classes Fault types
order jump call pred guard block def use
Project KLOC §2.2.1 §2.2.1 §2.2.1 §2.2.2 §2.2.2 §2.2.2 §2.3.1 §2.3.2 Avg per fault CF DF CF/DF All
Chart $96$ $1$ $3$ $5$ $4$ $11$ $3$ 1515{15}15 $7$ $1.88\pm 1.15$ $11$ $6$ $9$ $26$
Closure $90$ $6$ $9$ $37$ 4242{42}42 $53$ $30$ 8181{81}81 $6$ $1.98\pm 1.03$ $52$ $14$ $67$ $133$
Lang $22$ $2$ $18$ $11$ $11$ $31$ $11$ 4343{43}43 $8$ $2.08\pm 0.86$ $20$ $12$ $33$ $65$
Math $85$ $1$ $17$ $29$ $24$ $36$ $18$ 8383{83}83 $20$ $2.15\pm 1.04$ $22$ $27$ $57$ $106$
Mockito $11$ 00 $3$ $15$ $2$ $16$ $6$ 2828{28}28 $4$ $1.95\pm 0.86$ $10$ $4$ $24$ $38$
Time $28$ $3$ $9$ $6$ 77{7}7 $17$ $3$ 2121{21}21 $5$ $2.63\pm 1.61$ $6$ $4$ $17$ $27$
@@ 3. Dataset Construction Methodology
All $342$ $16$ $63$ $139$ $104$ $194$ $94$ 330330{330}330 $56$ $2.04\pm 1.03$ $154$ $81$ $253$ $488$
@@
We consider the difference in the flow graphs of the faulty and fixed code to manually assign fault classes according to the intrinsic fault nature. We infer the nature of the fault from the program context, taking into account the intended program semantics expressed by the repair. However, we investigate alternative repairs and compare all repair flow graphs to the faulty flow graph—the applied repair does not drive our classification.
Limitations
@@
We consider, similarly to Sobreira et al. (Sobreira et al., 2018), each Defects4J fault entry as a single fault that may display characteristics of multiple fault classes. However, certain entries do contain multiple independent faults (e.g., Chart-18). For comparability and compatibility with the repair classification (Sobreira et al., 2018), we do not distinguish between the multiple faults within an entry, and assign fault classes equally.
@@
Several of the Defects4J patches combine the repair with refactoring (see LABEL:listing:math46). This biases a repair classification and may inhibit our fault classification; we mitigate this by considering different possible repairs for a fault.
We consider only Defects4J projects, all of which are written in Java, as our benchmark set. Our results may not necessarily extend to projects of different types or programming languages. We leave this evaluation to future work.
4. Dataset Statistics
Figure 1 shows, for each individual project and aggregated over all projects, the distribution of the number of fault classes assigned to the faults. Most projects have a median (orange line) and upper quartile of two classes assigned to each fault. Overall, the majority of faults are each assigned between one and three fault classes, but some faults in Chart and Time are assigned as many as seven of the eight possible fault classes. All faults are assigned at least one fault class, as none of the faults were considered unclassifiable.
Figure 1. The distribution of number of fault classes assigned per fault, per project and overall